If the pressure of gas at 298 K is 1.23 atm, what would the temperature be if the pressure had increased to 2.25 atm


1
Expert's answer
2021-11-08T08:21:55-0500

Solution

In answering the question it is assumed that moles and volume are constant after temperature rise.


As the temperature is increased, pressure is increased if moles and volume remains constant.



"\therefore \>P \>\propto\>T"

"\implies \frac{P_1}{T_1}=\frac{P_2}{T_2}"


"P_1=1.23" "T_1=298"

"P_2=2.25" "T_2=T_f"


Substituting the value


"\frac{1.23}{298}=\frac{2.25}{T_f}"


Making "T_f" the subject ,


"T_f=\frac{2.25×298}{1.23}"


"=545 .12K"